When the pressure on a metal cube is increased by P Pa , its volume decreases to nV m3 .The bulk modulus of the material is (Initial volume is V m3)
 
(1)  2P/(1-n) Pa
(2) 2P/(1+n) Pa
(3) P/(1+n) Pa
(4) P/(1-n) Pa
 
Please help

Grade:12

2 Answers

Meet Shah
5 years ago
BULK MODULUS IS CHANGE IN PRESSURE PER RATE OF CHANGE IN VOLUME
 
BULK MODULUS= CHANGE IN PRESSURE * VOLUME
                                CHANGE IN VOLUME
=  /\frac{\Delta P\Delta V}{V}
\frac{PV}{V-nV}
=\frac{PV}{V(1-N))}
Meet Shah
5 years ago
 
BULK MODULUS IS CHANGE IN PRESSURE PER RATE OF CHANGE IN VOLUME
 
BULK MODULUS= \frac{CHANGE IN PRESSURE * VOLUME }{CHANGE IN VOLUME}
 
=\frac{\Delta P V}{\Delta V}
 
=\frac{P V}{V-nV}
 
=\frac{P}{1-n}
 
THEREFORE ANSWER (4) IS CORRECT